OpenStudy (anonymous):

is y-3*x-6=0 how you write y-3=3(x+1) in standard form * means multiply

OpenStudy (anonymous):

First off, standard form is y = mx + b where m is the slope of the line and b is the intercept. Now just solve for y. y - 3 = 3(x+1) y = 3(x+1) + 3 [add three to both sides] y = 3x + 3 + 3 [distribute 3 to (x+1)] y = 3x + 6 [add the 3s together to get 6]

OpenStudy (anonymous):

you taught me this better than my book your amazing thank you soooooo much

OpenStudy (anonymous):

Ha ha. Thanks. You make me think I should become a real tutor :)

OpenStudy (anonymous):

go for it I'm sure you'd be great